# Best Foot Forward: Robust Foot Reconstruction in-the-wild

### v1

@@ Abstract
Accurate 3D foot reconstruction is crucial for personalized orthotics, digital healthcare, and virtual fittings. However, existing methods struggle with incomplete scans and anatomical variations, particularly in self-scanning scenarios where user mobility is limited, making it difficult to capture areas like the arch and heel. We present a novel end-to-end pipeline that refines Structure-from-Motion (SfM) reconstruction. It first resolves scan alignment ambiguities using SE(3) canonicalization with a viewpoint prediction module, then completes missing geometry through an attention-based network trained on synthetically augmented point clouds. Our approach achieves state-of-the-art performance on reconstruction metrics while preserving clinically validated anatomical fidelity. By combining synthetic training data with learned geometric priors, we enable robust foot reconstruction under real-world capture conditions, unlocking new opportunities for mobile-based 3D scanning in healthcare and retail.
@@ 1 Introduction
Custom foot orthotics are essential for treating and preventing foot-related medical conditions by improving overall biomechanics [22]. Traditionally, they are manufactured using plaster casts and vacuum-forming, a costly and time-consuming process requiring in-person visits. Advances in digital scanning and additive manufacturing are transforming this field, enabling the creation of highly personalized orthotics that align with the principles of personalized medicine [12]. Beyond orthotics, high-fidelity 3D scanning has broader applications in both healthcare (e.g., custom prosthetics) and digital applications (e.g., virtual try-ons, gaming).
Despite advances in human body reconstruction [14, 10], the foot remains largely unexplored due to its complex biomechanics, high morphological variance, and imaging challenges like plantar surface occlusion. To address this, we present a novel, high-quality foot reconstruction method using multi-view mobile phone images, offering a robust and accessible solution for clinical and commercial use.
Building on advances in 3D computer vision, Structure-from-Motion (SfM) enables 3D reconstruction from 2D image sequences, while Multi-View Stereo (MVS) enhances geometric detail under controlled conditions [5]. However, self-scanning the foot remains challenging—users struggle to capture dense, overlapping views due to limited mobility and awkward angles, leading to incomplete image coverage (see Fig. 1). To address this, we make the following key contributions: (1) The first foot completion network to refine incomplete scans, improving robustness and accuracy in foot reconstruction. (2) A diverse foot dataset Hike3D with greater variation in attributes like age and height than previous datasets, enabling more robust modeling across different foot shapes. (3) Seamless integration with template-based foot reconstruction methods to generate high-quality meshes from partial point clouds. (4) A comprehensive evaluation showing our method outperforms COLMAP and state-of-the-art Gaussian splatting in robustness, feature accuracy, and surface quality.
@@ 2 Related Work
Early attempts in foot modeling relied on Principal Component Analysis (PCA) [1], but these models were simplistic, offering limited resolution and flexibility. Later approaches employ active sensor technologies, where structured light or depth cameras are used to generate point clouds [15, 28, 13]. However, the point cloud geometries obtained from these sensors are often noisy and incomplete. More recently, Boyne et al. proposed the FIND model [3] leveraging a template deformation strategy guided by an implicit neural network to improve reconstruction accuracy. Similarly, Osman et al. [18] developed SUPR, a PCA-based human foot model designed for seamless integration with the SMPL full-body model [14], enabling expressive and anatomically consistent reconstructions. However, both FIND and SUPR are limited by their training data, which strongly constrains the shape space. Our method, draws inspiration from multi-view reconstruction [23, 24, 8, 9] and shape completion [7, 25], both of which have proven effective in broader 3D reconstruction tasks. By leveraging these advancements, our approach can be seamlessly integrated into existing works, providing a more robust and generalizable solution for foot reconstruction.
@@ 3 Problem Setup
We consider a set of unposed images of the foot, denoted as =$\mathcal{I}=\{I_{1},I_{2},\dots,I_{N}\}$, where each image $I_{i}\in\mathbb{R}^{\mathrm{H}\times\mathrm{W}\times\mathrm{C}}$. Our objective is to reconstruct the complete geometry of the foot. To this end, we define a learnable function $\mathcal{F}_{\theta}$ that maps the image set $\mathcal{I}$ to a completed point cloud, such that $\mathcal{P}_{c}=\mathcal{F}_{\theta}(\mathcal{I}).$ To effectively address this, we decompose $\mathcal{F}_{\theta}$ into two composite functions $\mathcal{F}_{\theta}:=\mathcal{D}_{\theta}\circ\mathcal{S},$ where $\mathcal{S}:\mathcal{I}\to\mathbb{R}^{N\times 3}$ generates a dense, yet potentially incomplete, point cloud of the foot from the unposed images, and $\mathcal{D}_{\theta}:\mathbb{R}^{N\times 3}\to\mathbb{R}^{M\times 3}$ maps this partial point cloud to the completed point cloud target $\mathcal{P}_{c}\in\mathbb{R}^{M\times 3}$.
The challenge in learning $\mathcal{F}_{\theta}$ stems from supervision difficulties across inconsistent vector spaces. The geometric transformations between $\mathcal{D}_{\theta}$ and $\mathcal{S}$ remain unknown, creating constraints on pose and scale that complicate end-to-end system development. Our key insight addresses this by decomposing the problem into manageable sub-problems and leveraging synthetic training data at each stage. In the following section, we outline our method in more detail.
@@ 4 Method
We tackle complete foot reconstruction with a two-phase approach: first, we use Structure-from-Motion (SfM) and Multi-View Stereo (MVS) to estimate camera pose and generate an initial, though incomplete, point cloud; then, our shape completion module fills in missing geometry to produce a dense, complete representation. A naive approach to combining these two modules, estimating the geometric transform using iterative closest point (ICP) [2], often fails because the point clouds generated by SfM/MVS are typically incomplete. To overcome this, we introduce a viewpoint prediction (VPP) module, which provides a robust mechanism for estimating the transformation between the output of SfM/MVS and the expected input alignment for shape completion.
In the following section, we outline the core components of our reconstruction pipeline, illustrated in Fig. 2 (a). The pipeline begins with two branches: View-Point Prediction (VPP) and SfM & MVS, discussed in Sec. 4.1 and Sec. 4.2, respectively. The VPP module canonicalises the recovered partial point cloud (Sec. 4.3) before proceeding with foot completion and reconstruction (Sec. 4.4).
@@ 4.1 View-point Prediction
The first branch of our architecture, the VPP module, estimates both a bounding box of the foot and the pose relative a predefined template mesh. Given an unposed image set $\mathcal{I}$ and a reference mesh $\mathcal{M}_{\mathrm{ref}}$, we train a neural network to regress the approximate six degrees of freedom (6-DoF) of the camera pose relative to $\mathcal{M}_{\mathrm{ref}}$. Our method builds on YOLO6D [16], adopting a similar training strategy and leveraging synthetic data; implementation details are in Sec. 5. We represent the VPP module as $\mathcal{V}_{\phi}$ and define its output for a given image $\mathcal{I}_{i}\in\mathcal{I}$ as: $(\hat{\mathcal{C}}_{i},B_{i})=\mathcal{V}_{\phi}(\mathcal{I}_{i}),$ where $\hat{\mathcal{C}}_{i}$ denotes the estimated camera parameters, and $B_{i}$ represents the bounding box of the foot in the image.
@@ 4.2 SfM & MVS
We use a standard structure-from-motion (SfM) pipeline to estimate 3D structure by matching keypoints across views and jointly refining camera poses and a sparse point cloud via bundle adjustment. Specifically, we utilize GLOMAP [20], which from our experiments, observed to give significantly more efficient and scalable global reconstruction compared to COLMAP [23, 24]. For the image-set $\mathcal{I}$, we model the SfM process as $\boldsymbol{\mathcal{C}}=\mathrm{SfM}(\mathcal{I}),$ where each $\mathcal{C}_{i}\in\boldsymbol{\mathcal{C}}$ represents the estimated camera parameters of image $\mathcal{I}_{i}$. Using the bounding box $B_{i}$ from the VPP module, we generate segmentation masks via the Segment Anything Model 2 (SAM2) [21] in a zero-shot setting. Denoting the set of all bounding box centers as $\mathcal{B}$, we model the segmentation process as ^${\hat{\mathcal{I}}}=\mathrm{SAM}(\mathcal{I},\mathcal{B}),$ where ^$\hat{\mathcal{I}}_{i}\in{\hat{\mathcal{I}}}$ is the $i$-th masked image. To reconstruct a dense point cloud, we employ a multi-view stereo (MVS) approach [6], which estimates depth by matching pixel correspondences across multiple views and refining depth maps; in this work, We leverage the state-of-the-art MVSFormer++ [4] to recover high-quality point-clouds. Using the camera parameters from GLOMAP and segmentation masks from SAM2, we reconstruct the visible foot geometry pointcloud as $\mathcal{P}_{p}=\mathrm{MVS}(\boldsymbol{\mathcal{C}},{\hat{\mathcal{I}}}).$
Figure 2: (a) An overview of our reconstruction pipeline, more details in Sec. 4.
4.3 Point Cloud Canonicalisation
Partial point-clouds recovered from image sets have arbitrary poses and scales, complicating their use in a downstream shape completion module. To address this, we transform the point clouds into a known canonical frame using the camera parameters $\hat{C}_{i}$ estimated by the VPP module and the depth maps estimated in the MVS process $\mathcal{D}_{i}$; we present our canonicalisation in Algorithm. 1.
@@ 4.4 Point Cloud Completion
Building on recent attention-based approaches to point cloud modeling [26], we formulate completion as an auto-encoding problem, where the model predicts a global latent representation to guide the reconstruction. Our attention mechanism aggregates information across the entire point cloud, capturing both local details and global structural patterns without relying on predefined neighborhood structures. We adopt a coarse-to-fine reconstruction strategy with a scaffold-based skip connection that directly integrates a subset of the input point cloud into the reconstruction process. This scaffold helps maintain fidelity to the observed geometry while enabling the model to infer missing regions effectively.
In our standard pipeline, we then employ the screened Poisson surface reconstruction (SPSR) algorithm [9] to generate a mesh, using normals estimated via a $k$-nearest neighbors approach to ensure a smooth and consistent surface.
@@ 5 Implementation
Datasets: High-fidelity foot geometry datasets are scarce. Foot3D [3] is a valuable resource, but its narrow age and height range prompted us to develop Hike3D, a more diverse dataset for orthotics research. To broaden demographic coverage and strengthen design robustness, we integrate Hike3D with Foot3D. We release Hike3D as part of this work.
VPP module: We train the VPP module using synthetically generated images of meshes from our training set. To ensure diversity, we utilize 740 HDR backgrounds, creating various background combinations for our 50k synthetic images. We then fine-tune the model using 5k real images. Throughout the process, we apply the same augmentations and loss functions as in the original work [16].
Foot Completion Module: We train the foot completion module using a simulated scanning setup to generate paired partial and complete geometries. To improve robustness against noise and SE(3) perturbations, we apply data augmentations during training. Our dataset combines Hike3D and Foot3D, with a 1:8 training-to-test split. Each mesh underwent 10 spatial transformations (shifts, scaling, rotations), followed by five virtual scans per transformation, yielding 2000 training and 250 testing pairs. Supervision is applied by minimizing the Chamfer distance between predicted and ground-truth point clouds at intermediate steps of the network.

### v2

@@ Abstract
High-fidelity 3D foot reconstruction is crucial for prescription orthotics but is hindered by expensive, specialized equipment that limits patient access. We overcome this barrier with the first end-to-end pipeline to reconstruct clinically-accurate foot meshes from simple, self-captured smartphone videos. Our method uniquely solves the core challenges of in-the-wild scanning: we resolve pose ambiguities using SE(3) canonicalization with viewpoint prediction, and then complete partial geometry using an attention-based network. Clinical validation demonstrates that our reconstructions achieve state-of-the-art accuracy and meet prescription-readiness standards, preserving the anatomical fidelity essential for medical intervention. By democratizing high-quality foot assessment, our work unlocks new opportunities for accessible telemedicine, preventative diabetic care, and personalized orthotic treatment. We will release our dataset online at: https://bestfootforward.netlify.app.
@@ 1 Introduction
High-fidelity digital twins of the human body represent a foundational technology in autonomous healthcare, with the capacity to deliver significant advancements in personalized medicine. Toward this goal, we present a novel method for high-quality foot geometry capture, tailored to the design of custom foot orthotics. The proposed vision-based workflow automates the acquisition of complex foot geometries, demonstrating robustness against incomplete capture, to enable a fully digital manufacturing process consistent with the tenets of personalized medicine [18]. The methodology has direct implications for a spectrum of applications, including custom prosthetics, preoperative planning, and immersive digital environments such as virtual try-ons and gaming.
While significant progress has been made in reconstructing human anatomy from images [3, 21, 11, 14], creating accurate models of deformable and self-occluding parts like the human foot remains a formidable challenge. Self-captured scanning with commodity devices remains an unsolved problem while traditional approaches, including laser scanners [34], structured-light booths [7], marker-based gait laboratories [26], and impression boxes [19], rely on costly hardware, controlled environments, and skilled operators. Although recent efforts have sought to democratize high-quality foot reconstruction via smartphone video [17, 5], these approaches often depend on LIDAR-equipped devices, limiting accessibility, or they are not robust to in-the-wild partial scans, leading to unreliable geometry. To address this gap, we present the first method that robustly reconstructs high-quality foot geometry from a single, self-captured video taken with a standard smartphone.
While classical 3D reconstruction pipelines based on Structure-from-Motion (SfM) and Multi-View Stereo (MVS) excel under controlled conditions [8], they are notoriously brittle for in-the-wild medical capture. The self-scanning of a foot is a particularly challenging instance of this problem; user-captured sequences are often sparse and incomplete due to limited mobility (Fig. 1). Without a strong geometric prior, the resulting reconstructions are not suitable for designing effective orthotics. To overcome this fundamental limitation, we introduce a learning-based framework that robustly infers complete, high-fidelity foot geometry from imperfect data. Our primary contributions are:
  1. 1.
    A novel end-to-end pipeline for foot capture and completion, capable of accurately inferring the occluded geometry from partial and noisy point clouds.
  2. 2.
    The introduction of a comprehensive dataset of 3D foot scans (Hike3D), featuring broader demographic diversity than existing public datasets.
  3. 3.
    Extensive experiments demonstrating that our method significantly outperforms both classical pipelines (COLMAP) and state-of-the-art neural rendering techniques (e.g., 3D Gaussian Splatting) in reconstruction robustness, geometric accuracy, and orthotic design suitability.
@@ 2 Related Work
Multi-view reconstruction: Recovering 3D geometry from multiple images fundamentally depends on accurate estimation of camera intrinsics and relative poses. These parameters can be acquired through onboard sensors such as Inertial Measurement Units (IMUs) or estimated via sparse reconstruction techniques like Structure-from-Motion (SfM) [32]. Once the camera parameters are established, dense reconstruction is typically performed using Multi-View Stereo (MVS) [33], which computes consistent depth and normal maps across multiple views to produce an oriented point cloud. This point cloud is then converted into a surface mesh using a reconstruction method such as Poisson Surface Reconstruction [13]. A widely used implementation of this full pipeline is COLMAP [32, 33]. Recent advances have explored augmenting traditional multi-view reconstruction with learned priors, leveraging deep neural networks to improve accuracy and robustness. While these methods often incur higher computational costs and longer training times, they have demonstrated improved performance in challenging reconstruction scenarios. Neural rendering methods, such as Neural Radiance Fields (NeRF) [25] and 3D Gaussian Splatting [15], have also become dominant. While visually impressive, current approaches are typically slow to train, require many input views, and lack strong geometric priors, limiting their robustness for accurate 3D reconstruction. In our pipeline, we adopt MVSFormer++ [6], a transformer-based architecture designed to enhance multi-view stereo estimation through learned attention.
Pointcloud Shape Completion: Incomplete or sparse point clouds are a common limitation in 3D reconstruction pipelines, often resulting from occlusions, limited viewpoints, or sensor noise [30, 37]. Shape completion methods aim to infer missing geometry and recover plausible object or scene structures from partial observations [41]. Traditional approaches rely on geometric priors or optimization-based techniques [30, 10], while recent methods leverage deep learning to learn shape priors from large datasets [40, 36]. These learned models have demonstrated strong generalization capabilities, particularly in filling in large missing regions with semantically consistent geometry.
Human Foot Reconstruction: Early attempts in foot modeling relied on Principal Component Analysis (PCA) [1], but these models offered limited flexibility with restrictive shape spaces. Later approaches employ active sensor technologies, where structured light or depth cameras are used to generate point clouds [22, 39, 20]. However, the point cloud geometries obtained from these sensors are often noisy and incomplete. More recently, Boyne et al. proposed the FIND model [4] leveraging a template deformation strategy guided by an implicit neural network to improve reconstruction accuracy. Similarly, Osman et al. [27] developed SUPR, a PCA-based human foot model designed for seamless integration with the SMPL full-body model [21], enabling expressive and anatomically consistent reconstructions. However, neither FIND nor SUPR are well suited for handling partial or noisy inputs.
Figure 2: An overview of our reconstruction pipeline: Given an unposed image set $\mathcal{I}$, our method runs two parallel branches: SfM for camera calibration and point cloud generation, and pose estimation. We then canonicalize the point cloud with $\mathcal{T}\in\mathrm{SE}(3)$, apply shape completion $\mathcal{F}_{\theta}$, and use Poisson reconstruction for the final mesh.
@@ 3 Problem Setup
We consider a set of unposed images of the foot, denoted as $\mathcal{I}=\{I_{1},I_{2},\dots,I_{N}\}$, where each image $I_{i}\in\mathbb{R}^{\mathrm{H}\times\mathrm{W}\times\mathrm{C}}$. Our objective is to reconstruct the complete geometry of the foot. To this end, we define a learnable function $\mathcal{F}_{\theta}$ that maps the image set $\mathcal{I}$ to a completed point cloud, such that $\mathcal{P}_{c}=\mathcal{F}_{\theta}(\mathcal{I}).$ To effectively address this, we decompose $\mathcal{F}_{\theta}$ into two composite functions $\mathcal{F}_{\theta}:=\mathcal{D}_{\theta}\circ\mathcal{S},$ where $\mathcal{S}:\mathcal{I}\to\mathbb{R}^{Q\times 3}$ generates a dense, yet potentially incomplete, point cloud of the foot from the unposed images, and $\mathcal{D}_{\theta}:\mathbb{R}^{Q\times 3}\to\mathbb{R}^{M\times 3}$ maps this partial point cloud to the completed point cloud target $\mathcal{P}_{c}\in\mathbb{R}^{M\times 3}$.
The challenge in learning $\mathcal{F}_{\theta}$ stems from supervision difficulties across inconsistent vector spaces. The geometric transformations between $\mathcal{D}_{\theta}$ and $\mathcal{S}$ are in general unknown, creating constraints on pose and scale that complicate end-to-end system development. Our key insight is to address this by decomposing the problem into manageable sub-problems and leveraging synthetic training data at each stage. Then by leveraging a robust canonicalization approach we can effectively connect each stage together. In the following section, we outline our method in more detail.
@@ 4 Method
We tackle complete foot reconstruction with a two-phase approach: first, we use Structure-from-Motion (SfM) and Multi-View Stereo (MVS) to estimate camera pose and generate an initial, though potentially incomplete, point cloud. Then, our shape completion module completes missing geometry to produce a dense, complete representation.
A straightforward strategy is to estimate the geometric transformation between the two spaces using Iterative Closest Point (ICP) [2]. However, this often fails because the point clouds produced by SfM/MVS are typically sparse and incomplete, leading to unreliable alignment. To address this, we introduce a viewpoint prediction (VPP) module that robustly estimates the transformation between the SfM/MVS output and the expected input alignment for shape completion.
In the following section, we outline the core components of our reconstruction pipeline, illustrated in Fig. 2. The pipeline begins with two branches: View-Point Prediction (VPP) and SfM & MVS, discussed in Sec. 4.1 and Sec. 4.2, respectively. The VPP module canonicalizes the recovered partial point cloud (Sec. 4.3) before proceeding with foot completion and reconstruction (Sec. 4.4).
@@ 4.1 View-point Prediction
The first branch of our architecture, the VPP module, estimates both a bounding box of the foot and the pose relative to a predefined template mesh. Given an unposed image set $\mathcal{I}$ and a reference mesh $\mathcal{M}_{\mathrm{ref}}$, we train a neural network to regress the approximate six degrees of freedom (6-DoF) of the camera pose relative to $\mathcal{M}_{\mathrm{ref}}$. Our method builds on YOLO6D [23], adopting a similar training strategy and leveraging synthetic data; implementation details are in Sec. 5. We represent the VPP module as $\mathcal{V}_{\phi}$ and define its output for a given image $\mathcal{I}_{i}\in\mathcal{I}$ as: $(\hat{\mathcal{C}}_{i},B_{i})=\mathcal{V}_{\phi}(\mathcal{I}_{i}),$ where $\hat{\mathcal{C}}_{i}$ denotes the estimated camera parameters, and $B_{i}$ represents the bounding box of the foot in the image.
@@ 4.2 SfM & MVS
We use a standard structure-from-motion (SfM) pipeline to estimate 3D structure by matching keypoints across views and jointly refining camera poses and a sparse point cloud via bundle adjustment. Specifically, we utilize GLOMAP [29], which from our experiments, observed to give significantly more efficient and scalable global reconstruction compared to COLMAP [32, 33]. For the image-set $\mathcal{I}$, we model the SfM process as:
$\underbrace{\{\mathcal{C}_{1},\cdots,\mathcal{C}_{N}\}}_{\boldsymbol{\mathcal{C}}}=\mathrm{SfM}(\mathcal{I}),$
where each $\mathcal{C}_{i}\in\boldsymbol{\mathcal{C}}$ represents the estimated camera parameters of image $\mathcal{I}_{i}$. Using the bounding box $B_{i}$ from the VPP module, we generate segmentation masks via the Segment Anything Model 2 (SAM2) [31] in a zero-shot setting. Denoting the set of all bounding box centers as $\mathcal{B}$, we model the segmentation process as:
[…]
where $\hat{\mathcal{I}_{1}}$ is the $i$-th masked image. To reconstruct a dense point cloud, we employ a multi-view stereo (MVS) approach [9], which estimates depth by matching pixel correspondences across multiple views and refining depth maps; in this work we leverage the state-of-the-art MVSFormer++ [6] to recover a high-quality point-cloud. Using the camera parameters from GLOMAP and segmentation masks from SAM2, we reconstruct the visible foot geometry pointcloud as:
[…]
4.3 Point Cloud Canonicalization
Partial point-clouds recovered from image sets have arbitrary poses and scales, complicating their use in downstream shape completion. To address this, we transform the point clouds into a known canonical frame using the camera parameters $\hat{C}_{i}$ estimated by the VPP module and the depth maps estimated in the MVS process $\mathcal{D}_{i}$. We present our canonicalization approach in more detail in Algorithm. 1.
@@ 4.4 Point Cloud Completion
Building on recent attention-based approaches to point cloud modeling [35], we formulate completion as an auto-encoding problem, where the model predicts a global latent representation to guide the reconstruction. Our attention mechanism aggregates information across the entire point cloud, capturing both local details and global structural patterns without relying on predefined neighborhood structures. We adopt a coarse-to-fine reconstruction strategy with a scaffold-based skip connection that directly integrates a subset of the input point cloud into the reconstruction process. This scaffold helps maintain fidelity to the observed geometry while enabling the model to infer missing regions effectively and reduce scan noise.
In our standard pipeline, we then employ the screened Poisson surface reconstruction (SPSR) algorithm [12] to generate a mesh, with normals estimated via PCA.
Algorithm 1 Canonicalization
0: Reference mesh: $\mathcal{M}_{\text{ref}}$,          Camera parameters: $C_{i}$ [SfM], $\hat{C}_{i}$ [VPP],          Depth maps: $D_{i}$, Point-cloud: $\mathcal{P}_{p}$.
1: Select $k$ points $p_{1},p_{2},\ldots,p_{k}$ from $\mathcal{M}_{\text{ref}}$
2:for $i\leftarrow 1$ to NN do
3:  for $j\leftarrow 1$ to $k$ do
4:   $q_{i,j}\leftarrow\text{Project}(p_{j},\hat{C}_{i})$
5:   $d_{i,j}\leftarrow D_{i}(q_{i,j})$
6:   $p^{\prime}_{i,j}\leftarrow\text{BackProject}(q_{i,j},d_{i,j},C_{i})$
7:  end for
8:end for
9: Compute Centroids $c_{j}$ for each $j=1\ldots k$
10: Procrustes: $\{R,t,s\}\leftarrow\text{Procrustes}(\{p_{j}\},\{c_{j}\})$
[…]
12: ICP refinement: $\{R^{\prime},t^{\prime}\}\leftarrow\text{ICP}(\mathcal{P}^{\prime}_{p},\mathcal{M}_{\text{ref}})$
[…]
14:return  $\mathcal{P}_{\text{aligned}}$
@@ 5 Implementation
In this section we provide details of the implementation and training procedure used to construct our reconstruction pipeline.
Datasets: High-fidelity foot geometry datasets are scarce. Foot3D [4] is a valuable resource, but its narrow age and height range prompted us to develop Hike3D, a more diverse dataset for orthotics research. The Hike3D dataset comprises 15 participants recruited in the United States from a variety of occupational backgrounds. Data acquisition was performed using an EinStar structured-light scanner in a controlled indoor setting. Participants were seated with one foot extended, barefoot, in a neutral posture. Each foot was scanned individually, beginning from the plantar surface and sweeping around to the dorsal surface, capturing full geometry from bottom to top. Scans were cleaned, aligned to a common reference frame, and manually inspected for quality control, with their geometric accuracy evaluated in an offline assessment. To broaden demographic coverage and strengthen design robustness, we integrate Hike3D with Foot3D. We provide an overview of the distribution of Hike3D compared with [4] in Fig. 3. We will release the dataset as part of this work.
VPP module: We train the VPP module using synthetically generated images of meshes from our training set. To ensure diversity, we utilize 740 HDR backgrounds, creating various background combinations for our 50k synthetic images. We use Blender to simulate lighting variations and skin tone modulation, enhancing the realism of the synthetic data. We then fine-tune the model using 5k real images. Throughout the process, we apply the same augmentations and loss functions as in the original work [23].
Foot Completion Module: We train the foot completion module using a simulated scanning setup to generate paired partial and complete geometries. To improve robustness against noise and SE(3) perturbations, we apply data augmentations during training. Our dataset combines Hike3D and Foot3D, splitting the dataset with 80% for training and 20% for testing. Each mesh was augmented with 10 spatial transformations (shifts, scaling, rotations), followed by five virtual scans per transformation, yielding 2000 training and 250 testing pairs. Supervision is applied by minimizing the Chamfer distance between predicted and ground-truth point clouds at intermediate steps of the network.
Figure 3: Boxplots summarizing the distribution of heights and ages in our dataset, Hike3D, reveal a broader diversity in heights and significantly greater variation in ages compared to prior datasets. This age variation is particularly important, as feet undergo changes and deformations over time due to factors like footwear and activity levels.